OpenStudy (anonymous):

find the partial derivatives ∂z/∂y|(1, 1/2) if z=e^(x+2y) sin y

OpenStudy (turingtest):

where are you stuck here?

OpenStudy (anonymous):

\[z=e^{x+2y}\sin y\\ \text{By the product rule,}\\ \frac{\partial z}{\partial y}=\frac{\partial}{\partial y}\left[e^{x+2y}\right]\sin y+e^{x+2y}\frac{\partial}{\partial y}\left[\sin y\right]\]

OpenStudy (anonymous):

@SithsAndGiggles thank now i see how to solve it

OpenStudy (anonymous):

2e^x+2y sin + e^x+2y (cos y)

OpenStudy (anonymous):

@TuringTest sorry for the later reply i was afk for a while hopefully some1 alaready answer my quest